OpenStudy (anonymous):

Will give Medal!! What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 6 x - y = 3

OpenStudy (xapproachesinfinity):

first write y in terms of x in the second equa y=x+3 then replace y by x+3 in the first equa then solve for x

OpenStudy (anonymous):

Soooo x+3=x+3...? @xapproachesinfinity

OpenStudy (xapproachesinfinity):

no 4x+2(x-3)=6 4x+2x-6=6 6x=12 x=2

OpenStudy (anonymous):

Ohhh okay. Thank you I was confused.

OpenStudy (xapproachesinfinity):

sorry the y should be y=x-3 not x+3 a typo

OpenStudy (xapproachesinfinity):

you're welcome